OpenStudy (ashleyisakitty):

What is the equation of the function y=2/x translated 3 units to the left and 4 units up?

OpenStudy (johnweldon1993):

Same thing as a simpler function...to translate up or down, add or subtract a number from the whole function...to translate left or right, add or subtract a number from the 'x'

OpenStudy (ashleyisakitty):

y=2/x-3+4?

OpenStudy (johnweldon1993):

\[\large y = \frac{2}{x + 3}+4\] remember when you add a number to the 'x' it is making the opposite from what you expect to happen happen...so adding 3 would move to the LEFT as opposed to the right

OpenStudy (johnweldon1993):

IF you have a hard time visualizing that...think about this... Before you had just 'x' ...if you wanted 0...you would just make x = 0 Well if you add 3 to it...what does 'x' need to be to be 0 now? 0? no of course not....it would be -3 because -3 + 3 = 0 Hope that makes sense :)

OpenStudy (ashleyisakitty):

Yeah, I think i get it :) Thanks John :D

OpenStudy (johnweldon1993):

Of course :)